OpenStudy (anonymous):

Write an equation in slope intercept form for the following points and slope (4, -3), m = -1

TheSmartOne (thesmartone):

Slope intercept form is y=mx+b so we know the slope so the equation is y=-1x+b We know that the points (4,-3) work in this equation so plug it in. -3=-1(4)+b -3=-4+b b=1 So your final equation is y=-1x+1

TheSmartOne (thesmartone):

@tori1230